OpenStudy (anonymous):

how do you solve this system of equations? y= -x -2 y= 3x + 2

OpenStudy (anonymous):

If y=y then...?

OpenStudy (anonymous):

the subject is on solving systems of equations... i'm trying to find the ordered pair that goes with this..

OpenStudy (anonymous):

Exactly. So you are looking for an ordered pair (x,y) that satisfies both equations. So if you can solve for x you can plug it into one of the equations and get y.

OpenStudy (anonymous):

Multiply the top equation by 3 that way the x's will cancel out when you add them up. 3y = -3x - 6 y = 3x + 2 4y = -4 y = -1 Substitute back in -1 = 3x + 2 -3 = 3x -1 = x Answer (-1,-1)

OpenStudy (radar):

Or another approach, just add the equations as they stand. getting the sum as 2y=2x. Hey x = y now substitute whichever is easier for you. You will still get the same answer as blexting.